# Distributional Evaluation of Generative Models via Relative Density Ratio

### v1

@@ 2.3 Connection to existing methods
The main appeal of using $r(x)$ rather than the density ratio $g(x)$ is that, with a mixture of $p(x)$ and $q(x)$, the denominator measure is guaranteed to have full support over the numerator measure $p(x)$, i.e., $P\ll\frac{P+Q}{2}$, whereas the same is not guaranteed for the density ratio $g(x)$, especially in high-dimensional problems. Even though we could assume in theory fully overlapping support for $p$ and $q$; in practice, even for simple 1D Gaussians, the estimated density ratio $p/q$ can be far less than the theoretical value when $p$ has a nonzero measure but $q$ is near-zero. The theoretical properties related to $r(x)$ in Proposition 1 yield substantial improvements in numerical estimation compared to $g(x)$. Figure 1 shows the 1D normal comparison of $r(x)$ and $g(x)$ as the mean shift increases. We use a simple 4-layer MLP to learn both $r(x)$ and $g(x)$, and train the functional estimator on the observed training data with 1000 independent draws from both groups, and evaluate the trained function $r(x)$ or $g(x)$ over an evenly-spaced 1D grid of 500 grid points on [6,6][-6,6].
We observe that for both $\hat{g}(x)$ and $\hat{r}(x)$, the theoretical ratios can be larger than the estimated ones when $q$ is near-zero but $p$ is nonzero. There are two potential culprits. First, neural network functional estimators, such as a simple ReLU MLP, extrapolate linearly to out-of-sample areas (Xu et al., 2021) and are unable to adapt to nonlinear jumps in the theoretical ratio. Second, neural network estimators use a gradient-based approach. For ratio learning, when $f(x)=f_{1}(x)/f_{2}(x)$, given a loss function $\mathcal{L}\big(\hat{f}(x;\theta)\big)$, the gradient approach updates the neural network parameters $\theta$ by $\theta^{\text{new}}=\theta-\eta\nabla_{\theta}\mathcal{L}\big(\hat{f}(x;\theta)\big)$, where $\nabla_{\theta}\mathcal{L}(\hat{f})=\frac{\partial\mathcal{L}}{\partial\hat{f}}\,\frac{\partial\hat{f}}{\partial\theta}$ and $\frac{\partial\hat{f}}{\partial\theta}=\frac{1}{f_{2}}\frac{\partial f_{1}}{\partial\theta}-\frac{f_{1}}{f_{2}^{2}}\frac{\partial f_{2}}{\partial\theta}.$ It is more sensitive to the incremental changes in the denominator functions ($q$ or $\frac{p+q}{2}$) than to the numerator changes in $p$, and it tends to perform poorly when $p(x)$ has huge changes but $q(x)$ stays near-zero.
Figure 1: 1D normal comparison between the relative density ratio $r(x)$ and the density ratio $g(x)$. Both functional estimators are learned using the same ReLU MLP.
@@
Furthermore, when $p$ and $q$ are different, we need to measure how different they are. Here, we use the squared Hellinger distance as an example in Figure 1. Both the estimated $H^{2}(p,q)$ and $H^{2}(p,\frac{p+q}{2})$ increase as the two densities increasingly shift away. The theoretical upper bound for $H^{2}(p,q)$ is 1, while the upper bound for $H^{2}(p,\frac{p+q}{2})$ is $1-1/\sqrt{2}$ (shown in Proposition 1). In the far-right panel in Figure 1, when the normal means are shifted by 8, the empirical distributions are almost completely non-overlapping, which should yield the theoretical upper bound for both $H^{2}(p,q)$ and $H^{2}(p,\frac{p+q}{2})$. However, because of the massive extrapolation error induced by $\hat{g}(x)$, the estimated $\hat{H}^{2}(p,q)$ is still far from 1, whereas the estimated $\hat{H}^{2}\left(p,\frac{p+q}{2}\right)$ is almost exactly at the theoretical upper bound.
@@
then the density ratio can be recovered from the estimated $P(Z=1|X)$ through the “density ratio trick” because $g(X)=P(Z=1|X)/P(Z=0|X)\cdot P(Z=0)/P(Z=1)$ and $r=2g/(g+1)$. However, training a binary classifier under misclassification risk can be sensitive to the marginal sample probabilities $P(Z=i),i\in\left\{0,1\right\}$, and Awaya et al. (2025) demonstrate numerically that the density ratios estimated with binary classifiers can be severely biased when the two sample sizes are not equal and/or when the support of the two densities have regions of poor overlap. Estimation of $r(x)$ through the variational form of $\phi$-divergence is generally robust to unbalanced sample sizes and poor support overlap. This is highly relevant because both are common scenarios in applications involving generative models.
@@
A recent popular evaluation criterion for generative models is the precision-recall curve (Simon et al., 2019; Sajjadi et al., 2018). As illustrated in Figure 2, in the context of Simon et al. (2019), precision measures how much of $Q$ can be generated by a part of $P$, and recall measures how much of $P$ can be generated by a part of $Q$. The left panel in Figure 2 (partial precision, full recall) indicates that the generator has high fidelity (all $Q$-mass within $\text{supp}(P)$), but low coverage ($\text{supp}(Q)\subsetneq\text{supp}(P)$). The middle panel is the reversed case, where the generator has full coverage but low fidelity ($\text{supp}(P)\subsetneq\text{supp}(Q)$), and the right panel is a low-fidelity and low-coverage case ($\text{supp}(P)\nsubseteq\text{supp}(Q)$, $\text{supp}(Q)\nsubseteq\text{supp}(P)$). The second row in Figure 2 shows the theoretical and estimated $r(x)$, which gives a clear indication of the precision-recall/coverage-fidelity situation in each case; moreover, $r(x)$ pinpoints exactly which areas in the support suffer from low-fidelity (when $r(x)$ is near 0) and which areas suffer from low-coverage (when $r(x)$ is near 2). The bottom row of Figure 2 shows the histogram of $r(X_{i})$ evaluated on the training data, where $X_{i}$ is drawn from $P$ or $Q$. The histogram also shows a clear range of $r(x)$: for the full recall/fidelity and partial precision/coverage case (left), the range of $r(x)$ is strictly bounded below 2. The reversed case (middle) $r(x)$ is bounded above 0. More interestingly, in the case of partial precision and partial recall, we see a spike of mass of $r(X_{i}),X_{i}\sim P$ concentrated around 2, another spike of $r(X_{i}),X_{i}\sim Q$ concentrated around 0, and a smaller group of samples with overlapping $r(X_{i})$ from both groups in the middle. This histogram plot will be useful in high-dimensional evaluation to pinpoint regions of low fidelity or coverage and regions of good representation when it is expensive to create evenly spaced grids.
Figure 2: A 1D 3-mixture of Beta densities demonstrating the precision-recall comparison, simulation of Figure 1 in Simon et al. (2019). Left: partial precision, full recall. Middle: partial recall, full precision. Right: partial recall and precision, mode reweighting.
@@ 3 Theoretical Analysis
Based on the M-estimator theory (van der Vaart and Wellner, 1996), with the objective function being the variational form of the squared Hellinger distance (3), we first provide the general consistency and convergence rate results for estimators that satisfy a certain entropy bound of the function space in Section 3.1. Then we explore the sparse neural network function space with 1-Lipschitz activation functions such as ReLU and provide a convergence rate result when the true ratio function is assumed to be in the anisotropic Besov space. In this section, we use $p_{0},q_{0}$ to specifically denote the true density functions for distributions $P,Q$. Denote $g_{0}$ as the true density ratio $g_{0}=p_{0}/q_{0}$, and $r_{0}$ as the true relative density ratio $r_{0}=\frac{2p_{0}}{p_{0}+q_{0}}$.
@@ 3.1 General convergence theory
The following theorem establishes the almost sure consistency of the density ratio function as an M-estimator for the variational form of the squared Hellinger distance. The main proof techniques follow directly from Nguyen et al. (2010), with the adaptation to find a metric that can lower bound a functional difference induced by the Hellinger distance.
Let $\mathcal{G}$ be the general functional class representing the search space for the optimization problem in the variational form. As discussed in Section 2, the variational form of the squared Hellinger divergence in Table 1 can be written as
@@
The functional class $\mathcal{G}$ contains some $g$ such that $g=g_{0}$ almost surely.
@@
Assumption 2 is used for the consistency proof. Assumption 3 is a standard assumption in empirical process theory about the envelop functions, and Assumption 4 is the entropy condition for the general estimator space $\mathcal{G}$. Similar assumptions as Assumptions 2 to 4 are made in Nguyen et al. (2010).
@@
This proof of Theorem 2 follows the proof of Theorem 1 in Nguyen et al. (2010), whose theory is based on the Kullback-Leibler divergence, with adaptation to the squared Hellinger distance.
Note that Assumption 4 is a commonly used assumption when $\mathcal{G}^{1/2}-g_{0}^{1/2}$ is replaced by some general functional class $\mathcal{F}$. However, this assumption characterizes the size of the functional class in a probabilistic statement through empirical measures $P_{n},Q_{m}$. In the following Lemma 2, we show that a bounded bracketing number with the population metric $L^{1}(P)$ implies that the entropy number with the empirical metric $L^{1}(P_{n})$ is of the order $o_{P}(n)$.
A bracketing number of a functional class $\mathcal{F}$ is defined as follows: for a metric $\|\cdot\|$, given any small $\delta>0$, a $\delta$-bracket is defined as a pair of functions such that $l\leq f\leq u$ and $\|l-u\|\leq\delta$. The bracketing number $N_{[]}(\delta,\mathcal{F},\|\cdot\|)$ is defined as the minimum number of $\delta$-brackets needed to cover $\mathcal{F}$.
@@
For a general functional class $\mathcal{F}$ that satisfies for any $\epsilon>0$, $N_{[]}(\epsilon,\mathcal{F},L^{1}(P))<\infty$ implies $\frac{1}{n}\log N(\epsilon,\mathcal{F},L^{1}(P_{n}))\overset{P}{\to}0$.
@@
Define the functional squared Hellinger distance with respect to measure $Q$ as
@@
We follow the notations in van der Vaart and Wellner (1996) and denote the bracketing entropy integral as follows:
@@
Assumption 5 is stronger than Assumption 3. A similar assumption is also required for the convergence rate in Nguyen et al. (2010).
@@
Under Assumptions 1, 2, and 4 to 6, when replacing $Q$ by $\widetilde{Q}$, we have that $h_{\widetilde{Q}}(\hat{r}_{n},r_{0})=O_{p}(n^{-\frac{1}{2+\gamma}})$ where $\gamma=\min\left\{\gamma_{p}.\gamma_{q}\right\}$.
@@
3.2 Sieved M-estimator for Neural Networks
The above section provides convergence rate results for general estimators, as long as the estimator’s function space satisfies the entropy conditions in Assumption 6. In this section, we provide a convergence rate result in terms of sparse neural networks and assume that the true density $g_{0}$ is in a very general functional class, specifically the anisotropic Besov space (Nikol’skii, 2012; Triebel, 2006; Suzuki and Nitanda, 2021). Our result is built on Suzuki and Nitanda (2021), which provides the approximation error of the sparse neural network for functions in the anisotropic Besov space. In this section, we assume that the support of the density ratio is on $\mathcal{X}=[0,1]^{d}$.
@@ 3.2.1 A Brief Review on the Anisotropic Besov Space
We first introduce the anisotropic Besov space, following the same notations as in Suzuki and Nitanda (2021).
@@
To understand this, the anisotropic Besov space $B^{\beta}_{s,t}(\mathcal{X})$ is controlled by $\beta$, the smoothness vector that allows for different smoothness $\beta_{i}$ in each direction; $s$, indicating that the smoothness is measured in $L^{s}(\mathcal{X})$; and $t$, used in the seminorm (10). The seminorm (10) measures the localized anisotropic smoothness of $f$ in terms of the scale of oscillation. Unlike Sobolev spaces that measure smoothness through derivatives, the modulus of smoothness in (9) measures the finite difference given an incremental change $h$ and does not require differentiability. The seminorm (10) dyadically partitions the incremental change $h$ in different directions, $|h_{i}|<2^{-k/\beta_{i}}$, and $2^{k}$ is a normalization factor. Summing over $k$ aggregates the smoothness across different partition layers. Triebel (2011) and Suzuki and Nitanda (2021) provide more detailed results on the relation between Hölder space and anisotropic Besov space. For example (Proposition 1 in Suzuki and Nitanda (2021)), when $\beta=(\beta_{0},\dots,\beta_{0})\in\mathbb{R}^{d}$, and $\beta_{0}\notin\mathbb{N}$, then Hölder space $\mathcal{C}^{\beta_{0}}=B^{\beta}_{\infty,\infty}$.
The main advantage of assuming the true density ratio $g_{0}$ lies in $B^{\beta}_{s,t}(\mathcal{X})$ is that $B^{\beta}_{s,t}(\mathcal{X})$ allows for different smoothness in each direction. This coincides with our belief in two-sample comparisons in high dimensions, especially when comparing generative models: the two-sample difference may only lie in a few directions.

### v2

@@ 2.3 Connection to existing methods
The main appeal of using $r(x)$ rather than the density ratio $g(x)$ is that, with a mixture of $p(x)$ and $q(x)$, the denominator measure is guaranteed to have full support over the numerator measure $p(x)$, i.e., $P\ll\frac{P+Q}{2}$, whereas the same is not guaranteed for the density ratio $g(x)$, especially in high-dimensional problems. Even though we could assume in theory fully overlapping support for $p$ and $q$; in practice, even for simple 1D Gaussians, the estimated density ratio $p/q$ can be far less than the theoretical value when $p$ has a nonzero measure but $q$ is near-zero. The theoretical properties related to $r(x)$ in Proposition 1 yield substantial improvements in numerical estimation compared to $g(x)$. Figure 1 shows the 1D normal comparison of $r(x)$ and $g(x)$ as the mean shift increases. We use a simple 4-layer MLP to learn both $r(x)$ and $g(x)$, and train the function estimator on the observed training data with 1000 independent draws from both groups, and evaluate the trained function $r(x)$ or $g(x)$ over an evenly-spaced 1D grid of 500 grid points on [6,6][-6,6].
We observe that for both $\hat{g}(x)$ and $\hat{r}(x)$, there exist estimation errors. There are two potential culprits. First, neural network estimators, such as a simple ReLU MLP, extrapolate linearly to out-of-sample areas (xu2021how), which leads to estimation error in the low data mass region (see Figure 1 Column 1, $\hat{r}(x)$ has large extrapolation error when $x>3$). Second, the estimation performance of $r(x)$ depends on the output activation function of the last layer. There are different choices of the output activation function that all have a bounded image $[0,2]$, and the estimation performance depends on how fast the output activation function saturates to the end points $0,2$. In the Supplementary Section D, we provide a discussion on the choice of output activation functions and a sensitivity analysis on their performances.
Figure 1: 1D normal comparison between the relative density ratio $r(x)$ (Row 2) and the density ratio $g(x)$ (Row 3). Both function estimators are learned using the same ReLU MLP.
@@
Recall desideratum (A), when $p$ and $q$ are different, we need to measure how different they are. The objective value of (4) automatically provides the estimate for the squared Hellinger distance between $p$ and $(p+q)/2$. Hence, we use the squared Hellinger distance as the global numeric score in desideratum (A), as illustrated in Figure 1. Both the estimated $H^{2}(p,q)$ and $H^{2}(p,\frac{p+q}{2})$ increase as the two densities increasingly shift away. The theoretical upper bound for $H^{2}(p,q)$ is 1, while the upper bound for $H^{2}(p,\frac{p+q}{2})$ is $1-1/\sqrt{2}$ (shown in Proposition 1). In the far-right panel in Figure 1, when the normal means are shifted by 8, the empirical distributions are almost completely non-overlapping, which should yield the theoretical upper bound for both $H^{2}(p,q)$ and $H^{2}(p,\frac{p+q}{2})$. However, because of the massive extrapolation error induced by $\hat{g}(x)$, the estimated $\hat{H}^{2}(p,q)$ is still far from 1, whereas the estimated $\hat{H}^{2}\left(p,\frac{p+q}{2}\right)$ is almost exactly at the theoretical upper bound.
@@
then the density ratio can be recovered from the estimated $P(Z=1|X)$ through the “density ratio trick” because $g(X)=P(Z=1|X)/P(Z=0|X)\cdot P(Z=0)/P(Z=1)$ and $r=2g/(g+1)$. However, training a binary classifier under misclassification risk can be sensitive to the marginal sample probabilities $P(Z=i),i\in\left\{0,1\right\}$, and awaya2025two demonstrate numerically that the density ratios estimated with binary classifiers can be severely biased when the two sample sizes are not equal and/or when the support of the two densities have regions of poor overlap. Estimation of $r(x)$ through the variational form of $\phi$-divergence is generally robust to unbalanced sample sizes and poor support overlap. This is highly relevant because both are common scenarios in applications involving generative models.
@@
A recent popular evaluation criterion for generative models is the precision-recall curve (pmlr-v97-simon19a; sajjadi2018assessing). As illustrated in Figure 2, in the context of Figure 1 in pmlr-v97-simon19a, precision measures how much of $Q$ can be generated by a part of $P$, and recall measures how much of $P$ can be generated by a part of $Q$. The left panel in Figure 2 (partial precision, full recall) indicates that the generator has full coverage (all $P$-mass within $\text{supp}(Q)$), but suffers from potential model hallucination ($\text{supp}(Q)\subsetneq\text{supp}(P)$). The middle panel is the reversed case, where the generator has high fidelity but only partial coverage ($\text{supp}(Q)\subsetneq\text{supp}(P)$), missing critical regions in the real data, and the right panel is a low-fidelity and low-coverage case ($\text{supp}(P)\nsubseteq\text{supp}(Q)$, $\text{supp}(Q)\nsubseteq\text{supp}(P)$). The second row in Figure 2 shows the theoretical and estimated $r(x)$, which gives a clear indication of the precision-recall/coverage-fidelity situation in each case; moreover, $r(x)$ pinpoints exactly which areas in the support suffer from low-fidelity (when $r(x)$ is near 0) and which areas suffer from low-coverage (when $r(x)$ is near 2). The bottom row of Figure 2 shows the histogram of $r(X_{i})$ evaluated on the training data, where $X_{i}$ is drawn from $P$ or $Q$. The histogram also shows a clear range of $r(x)$: for the model hallucination (low fidelity) case (left), the range of $r(x)$ is strictly bounded below 2 with a large mass near 0. In the reversed case (middle), $r(x)$ is bounded above 0 with large mass near 2. More interestingly, in the case of partial precision and partial recall (right), we see a spike of mass of $r(X_{i}),X_{i}\sim P$ concentrated around 2, another spike of $r(X_{i}),X_{i}\sim Q$ concentrated around 0, and a smaller group of samples with overlapping $r(X_{i})$ from both groups in the middle. This histogram plot will be useful in high-dimensional evaluation to pinpoint regions of low fidelity or coverage and regions of good representation when it is expensive to create evenly spaced grids.
Figure 2: A 1D 3-mixture of Beta densities demonstrating the precision-recall comparison, simulation of Figure 1 in pmlr-v97-simon19a. Left (model hallucination, low fidelity): partial precision, full recall. Middle (low coverage): partial recall, full precision. Right: partial recall and precision, a mixed case.
@@ 3 Theoretical Analysis
Based on the M-estimator theory (van1996weak), with the objective function being the variational form of the squared Hellinger distance (3), we first provide the general consistency and convergence rate results for estimators that satisfy a certain entropy bound of the function space in Section 3.1. Then we explore the sparse neural network function space with 1-Lipschitz activation functions such as ReLU and provide a convergence rate result when the true ratio function is assumed to be in the anisotropic Besov space. In this section, we use $p_{0},q_{0}$ to specifically denote the true density functions for distributions $P,Q$. Denote $g_{0}$ as the true density ratio $g_{0}=p_{0}/q_{0}$, and $r_{0}$ as the true relative density ratio $r_{0}=\frac{2p_{0}}{p_{0}+q_{0}}$.
@@ 3.1 General convergence theory
The following theorem establishes the almost sure consistency of the density ratio function as an M-estimator for the variational form of the squared Hellinger distance. The main proof techniques follow directly from Nguyen2010-xb, with the adaptation to find a metric that can lower bound the difference between functions induced by the Hellinger distance.
Let $\mathcal{G}$ be the general function class representing the search space for the optimization problem in the variational form. As discussed in Section 2, the variational form of the squared Hellinger divergence in Table 1 can be written as
@@
The function class $\mathcal{G}$ contains some $g$ such that $g=g_{0}$ almost surely.
@@
Assumption 2 is used for the consistency proof. Assumption 3 is a standard assumption in empirical process theory about the envelop functions, and Assumption 4 is the entropy condition for the general estimator space $\mathcal{G}$. Similar assumptions as Assumptions 2 to 4 are made in Nguyen2010-xb.
@@
This proof of Theorem 2 follows the proof of Theorem 1 in Nguyen2010-xb, whose theory is based on the Kullback-Leibler divergence, with adaptation to the squared Hellinger distance.
Note that Assumption 4 is a commonly used assumption when $\mathcal{G}^{1/2}-g_{0}^{1/2}$ is replaced by some general function class $\mathcal{F}$. However, this assumption characterizes the size of the function class in a probabilistic statement through empirical measures $P_{n},Q_{m}$. In the following Lemma 2, we show that a bounded bracketing number with the population metric $L^{1}(P)$ implies that the entropy number with the empirical metric $L^{1}(P_{n})$ is of the order $o_{P}(n)$.
A bracketing number of a function class $\mathcal{F}$ is defined as follows: for a metric $\|\cdot\|$, given any small $\delta>0$, a $\delta$-bracket is defined as a pair of functions such that $l\leq f\leq u$ and $\|l-u\|\leq\delta$. The bracketing number $N_{[]}(\delta,\mathcal{F},\|\cdot\|)$ is defined as the minimum number of $\delta$-brackets needed to cover $\mathcal{F}$.
@@
For a general function class $\mathcal{F}$ and for any $\epsilon>0$, $N_{[]}(\epsilon,\mathcal{F},L^{1}(P))<\infty$ implies $\frac{1}{n}\log N(\epsilon,\mathcal{F},L^{1}(P_{n}))\overset{P}{\to}0$.
@@
Define the squared Hellinger metric for functions with respect to measure $Q$ as
@@
We follow the notations in van1996weak and denote the bracketing entropy integral as follows:
@@
Assumption 5 is stronger than Assumption 3. A similar assumption is also required for the convergence rate in Nguyen2010-xb.
@@
Under Assumptions 1, 2, and 4 to 6, when replacing $Q$ by $\widetilde{Q}$, we have $h_{\widetilde{Q}}(\hat{r}_{n},r_{0})=O_{p}(n^{-\frac{1}{2+\gamma}})$ where $\gamma=\min\left\{\gamma_{p}.\gamma_{q}\right\}$.
@@
3.2 Sieved Neural Network-based M-estimators
The above section provides convergence rate results for general estimators, as long as the function space of the estimator satisfies the entropy conditions in Assumption 6. In this section, we provide a convergence rate result in terms of sparse neural networks and assume that the true density $g_{0}$ is in a very general function class, specifically the anisotropic Besov space (nikol2012approximation; triebel2006theory; suzuki2021deep). Our result is built on suzuki2021deep, which provides the approximation error of the sparse neural network for functions in the anisotropic Besov space. In this section, we assume that the support of the density ratio is on $\mathcal{X}=[0,1]^{d}$.
@@ 3.2.1 A Brief Review on the Anisotropic Besov Space
We first introduce the anisotropic Besov space, following the same notations as in suzuki2021deep.
@@
To understand this, the anisotropic Besov space $B^{\beta}_{s,t}(\mathcal{X})$ is controlled by $\beta$, the smoothness vector that allows for different smoothness $\beta_{i}$ in each direction; $s$, indicating that the smoothness is measured in $L^{s}(\mathcal{X})$; and $t$, used in the seminorm (10). The seminorm (10) measures the localized anisotropic smoothness of $f$ in terms of the scale of oscillation. Unlike Sobolev spaces that measure smoothness through derivatives, the modulus of smoothness in (9) measures the finite difference given an incremental change $h$ and does not require differentiability. The seminorm (10) dyadically partitions the incremental change $h$ in different directions, $|h_{i}|<2^{-k/\beta_{i}}$, and $2^{k}$ is a normalization factor. Summing over $k$ aggregates the smoothness across different partition layers. triebel2011entropy and suzuki2021deep provide more detailed results on the relation between Hölder space and anisotropic Besov space. For example (Proposition 1 in suzuki2021deep), when $\beta=(\beta_{0},\dots,\beta_{0})\in\mathbb{R}^{d}$, and $\beta_{0}\notin\mathbb{N}$, then Hölder space $\mathcal{C}^{\beta_{0}}=B^{\beta}_{\infty,\infty}$.
The main advantage of assuming the true density ratio $g_{0}$ lies in $B^{\beta}_{s,t}(\mathcal{X})$ is that $B^{\beta}_{s,t}(\mathcal{X})$ allows for different smoothness in each direction. This coincides with our belief in two-sample comparisons in high dimensions, especially when comparing generative models: the two-sample difference may only lie in a few directions, especially when the generative models can produce realistic samples.